In a white water relay race, one kayak team averaged 26 km/h downstream but only 6 km/h upstream in the white water course. What was the rate of the kayak in still water? What was the rate of the current?
Answer by ankor@dixie-net.com(15652)

n a white water relay race, one kayak team averaged 26 km/h downstream but only 6 km/h upstream in the white water course.
What was the rate of the kayak in still water?
:
Let s = kayak speed in still water
Let c = rate of the current
:
From the information given, we know:
downstream: s + c = 26
and
upstream: s - c = 6
:
We can use elimination here:
s + c = 26
s - c = 6
-----------adding eliminates c, find s
2s = 32
s = 16 km/hr in still water
:
"What was the rate of the current?"
16 + c = 26
c = 26 - 16
c = 10 km/hr is the rate of the current.